# Interpreting Graph Inference with Skyline Explanations

### v2

@@ VII Conclusion
We have proposed a class of skyline explanations that simultaneously optimize multiple explainability measures for interpreting the output of graph neural networks. We have shown that the generation of skyline explanations (as an explanatory query processing problem) is nontrivial even for polynomially bounded input space and three measures. We have introduced feasible algorithms, sequential and parallel, for the skyline explanatory query processing, as well as its diversified counterpart, with provable quality guarantees in terms of Pareto optimality. Our experimental study has verified that our methods are practical for large-scale graphs and $\mathsf{GNN}$-based classification and generate more comprehensive explanations, compared with state-of-the-art $\mathsf{GNN}$ explainers.
@@ VII-A Details of Algorithms and Procedures
Procedure updateSX: Update Dominance Relation. We introduce the details of Procedure updateSX; specifically, we describe how the (1+$\epsilon$) dominance relations of exploratory subgraphs are dynamically maintained. With the stream of upcoming states, i.e., explanatory subgraphs, we maintain a lattice structure for the dominance relationships. Specifically, we first compare each newly arrived state $s$ with the current skylines, and incrementally update the dominance lattice by conducting one of the following steps:
@@
    If certain skylines (1+$\epsilon$)-dominate it, it inserts a directed edge from each dominating skylines to the new state; the checking of (1+$\epsilon$)-dominance takes constant time for small measure space involving several explainability measures;
@@
3:set ${\mathcal{G}}_{\epsilon}$:$\emptyset$;
4:identify edges for each hop: ${\mathcal{E}}$:=$\{E_{L},E_{L-1},\ldots,E_{1}\}$;
5:for $l$ = $L$ to $1$ do
6:     initializes state $s_{0}$:=$G^{l}(v_{t})$
@@
13:              ${\mathcal{G}}_{\epsilon}$:=updateDSX $(s,G_{s},\zeta^{\prime},{\mathcal{G}}_{\epsilon})$; $E_{l}$=$E_{l}\backslash\{e\}$;
@@
3:initializes state $s$ with structure $\Phi(G_{s})$:$\emptyset$;
@@
7:     $\Delta(s|{\mathcal{G}}_{\epsilon})$ := ${\sf DivS}($${\mathcal{G}}_{\epsilon}$$\cup$$\{G_{s}\}$$)-{\sf DivS}($${\mathcal{G}}_{\epsilon}$))))
8:     if $|{\mathcal{G}}_{\epsilon}|$$<$$k$ and $\Delta(s|{\mathcal{G}}_{\epsilon})$$\geq$$\frac{(1+\epsilon)/2-{\sf DivS}({\mathcal{G}}_{\epsilon})}{k-|{\mathcal{G}}_{\epsilon}|}$ then
@@
3:${\mathcal{G}_{Q}}$ := $\emptyset$; $V_{T}$ := $\bigcup_{i=1}^{n}{\mathsf{SXQ}}^{k}_{i}.{v_{t}}$; node partition $P$ := $\emptyset$;
@@
Parallel Algorithms. We first present the details of the baseline parallel algorithms ($\mathsf{ParaSX}$-$\mathsf{N}$ and $\mathsf{ParaSX}$-$\mathsf{EIS}$). They differ in modeling and minimizing the makespan of skyline explanatory query processing, i.e., the total elapsed (wall-clock) time from the start of the first ${\mathsf{SXQ}}^{k}$ to the completion of the last ${\mathsf{SXQ}}^{k}$ in a batch. Our parallel algorithms aim to tackle the common makespan minimization problem. Given a batch of ${\mathsf{SXQ}}^{k}$ queries $Q$ = $\{{\mathsf{SXQ}}^{k}_{1},{\mathsf{SXQ}}^{k}_{2},\cdots,{\mathsf{SXQ}}^{k}_{n}\}$ and a set of threads running in parallel: $T$ = $\{t_{1},t_{2},\cdots,t_{m}\}$, the objective is to assign each thread $t_{i}$ a subset of query workload $Q_{i}\subseteq Q$, such that the maximum time of any thread takes to complete its subset of queries (or its makespan) is minimized. More formally, it computes an assignment $\Psi$ : $Q$rightarrow $T$ to minimize
@@
Figure 8: Running Example for $\mathsf{ParaSX}$. The blue edges are computed when processing $v_{1}$; The red edge is computed when processing $v_{3}$. When a query batch containing six nodes arrives, $\mathsf{ParaSX}$ first partitions the queries into threads by maximizing intra-cluster Jaccard similarity. Within each thread, the queries are then ordered using the $k$-core algorithm. Consider a thread where the query order is $v_{1}$ followed by $v_{3}$. $\mathsf{ASX}$-$\mathsf{OP}$ initializes the shared edge information table by computing the edge scores (depicted in blue) from the neighborhood graph of $v_{1}$. During the processing of $v_{3}$, only one additional edge—$e_{1112}$ (shown in red)—requires computation, as the remaining edges are shared with $v_{1}$ and can directly reuse the values stored in the shared table.
@@ VII-B Additional Experimental Results
Impact of $k$. Setting ${\mathcal{M}}$ as $\mathsf{GCN}$-based classifier with $3$ layers, we vary $k$ from 1 to 25. Since our competitors are not configurable w.r.t. $k$, we show the IPF and IGD scores of their solutions (that are independent of $k$), along with our $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ (which are dependent on $k$) in Figures 7(a) and 7(b). As $k$ becomes larger, all achieve better scores and can consistently generate explanations with higher quality. This is because larger $k$ allows more dominating subgraphs to be verified, whenever possible for at least one measure. In contrast, our competitors are not aware of the growth of $k$, given that they may “stuck” at local optimal explanations that are optimized over a single measure, hence fail to improve the solution even with larger $k$. Moreover, $\mathsf{MOExp}$ does not explicitly constrain the size of the explanation set, which can result up to 200 explanatory subgraphs based on our empirical results. Meanwhile, even with $k$=$1$, our methods outperform other competitors for both IPF and IGD.
Varying $\mathsf{GNN}$ classes. Fixing $k$=$10$, we report IPF scores of $3$-layer $\mathsf{GNN}$ explainers for representative $\mathsf{GNN}$ classes including $\mathsf{GCN}$, $\mathsf{GAT}$, and $\mathsf{GIN}$. As verified in Figure 7(c), $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ consistently outperform other explanations for different $\mathsf{GNN}$ classes. This verifies that our approaches remain stable in generating high-quality explanations for different types of $\mathsf{GNNs}$. Our observations for IGD scores are consistent, hence omitted.
Varying number of layers. Fixing $k$=$10$, we report IGD scores of $\mathsf{GCN}$ explainers, with the number of layer $L$ varies from 1 to 3. over $\mathsf{Cora}$. Fig 7(d) shows that $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ are able to achieve better IGD scores for generating explanations with more complex $\mathsf{GNNs}$ having “deeper” architecture. Indeed, as $L$ becomes larger, our all three methods are able to identify more connected subgraphs that contribute to better ${\mathsf{fdl}}^{+}$ and ${\mathsf{fdl}}^{-}$ scores, hence improving explanation quality.

### v3

@@ VII Conclusion
We proposed a class of skyline explanations that optimize multiple explainability measures for interpreting GNN predictions. We showed that generating skyline explanations, viewed as an explanatory query processing problem, is nontrivial even with polynomially bounded input space and three measures. To address this, we developed sequential, parallel, and diversified algorithms with provable Pareto-optimality guarantees. Extensive experiments demonstrate that our methods are practical for large-scale graphs and GNN-based classification, producing more comprehensive explanations than state-of-the-art GNN explainers.
@@ VII-A Details of Algorithms and Procedures
[…]
@@
[…]
@@
3:set ${\mathcal{G}}_{\epsilon}$$\emptyset$;
4:identify edges for each hop: $\mathcal{E}$$\{E_{L},E_{L-1},\ldots,E_{1}\}$;
5:for $l$ = $L$ to 11 do
6:  initializes state $s_{0}$$G^{l}(v_{t})$
@@
[…]
@@
3:initializes state $s$ with structure $\Phi(G_{s})$$\emptyset$;
@@
7:  $\Delta(s|{\mathcal{G}}_{\epsilon})$ := ${\sf DivS}($${\mathcal{G}}_{\epsilon}\cup$$\{G_{s}\}$$)-{\sf DivS}($${\mathcal{G}}_{\epsilon}$))
8:  if $|{\mathcal{G}}_{\epsilon}|$<<$k$ and $\Delta(s|{\mathcal{G}}_{\epsilon})\geq$$\frac{(1+\epsilon)/2-{\sf DivS}({\mathcal{G}}_{\epsilon})}{k-|{\mathcal{G}}_{\epsilon}|}$ then
@@
[…]
@@
Parallel Algorithms. We first present the details of the baseline parallel algorithms ($\mathsf{ParaSX}$-$\mathsf{N}$ and $\mathsf{ParaSX}$-$\mathsf{EIS}$). They differ in modeling and minimizing the makespan of skyline explanatory query processing, i.e., the total elapsed (wall-clock) time from the start of the first ${\mathsf{SXQ}}^{k}$ to the completion of the last ${\mathsf{SXQ}}^{k}$ in a batch. Our parallel algorithms aim to tackle the common makespan minimization problem. Given a batch of ${\mathsf{SXQ}}^{k}$ queries $Q$ = $\{{\mathsf{SXQ}}^{k}_{1},{\mathsf{SXQ}}^{k}_{2},\cdots,{\mathsf{SXQ}}^{k}_{n}\}$ and a set of threads running in parallel: $T$ = $\{t_{1},t_{2},\cdots,t_{m}\}$, the objective is to assign each thread $t_{i}$ a subset of query workload $Q_{i}\subseteq Q$, such that the maximum time of any thread takes to complete its subset of queries (or its makespan) is minimized. More formally, it computes an assignment $\Psi$ : $Q$ $\rightarrow$ $T$ to minimize
@@
[…]
@@ VII-B Additional Experimental Results
Impact of $k$. Setting $\mathcal{M}$} as $\mathsf{GCN}$-based classifier with $3$ layers, we vary $k$ from 1 to 25. Since our competitors are not configurable w.r.t. $k$, we show the IPF and IGD scores of their solutions (that are independent of $k$), along with our $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ (which are dependent on $k$) in Figures 7(a) and 7(b). As $k$ becomes larger, all achieve better scores and can consistently generate explanations with higher quality. This is because larger $k$ allows more dominating subgraphs to be verified, whenever possible for at least one measure. In contrast, our competitors are not aware of the growth of $k$, given that they may “stuck” at local optimal explanations that are optimized over a single measure, hence fail to improve the solution even with larger $k$. Moreover, $\mathsf{MOExp}$ does not explicitly constrain the size of the explanation set, which can result up to 200 explanatory subgraphs based on our empirical results. Meanwhile, even with kk=11, our methods outperform other competitors for both IPF and IGD.
Varying $\mathsf{GNN}$ classes. Fixing kk=$10$, we report IPF scores of $3$-layer $\mathsf{GNN}$ explainers for representative $\mathsf{GNN}$ classes including $\mathsf{GCN}$, $\mathsf{GAT}$, and $\mathsf{GIN}$. As verified in Figure 7(c), $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ consistently outperform other explanations for different $\mathsf{GNN}$ classes. This verifies that our approaches remain stable in generating high-quality explanations for different types of $\mathsf{GNNs}$. Our observations for IGD scores are consistent, hence omitted.
Varying number of layers. Fixing kk=$10$, we report IGD scores of $\mathsf{GCN}$ explainers, with the number of layer $L$ varies from 1 to 3. over $\mathsf{Cora}$. Fig 7(d) shows that $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ are able to achieve better IGD scores for generating explanations with more complex $\mathsf{GNNs}$ having “deeper” architecture. Indeed, as $L$ becomes larger, our all three methods are able to identify more connected subgraphs that contribute to better $\mathsf{fdl}$$^{+}$ and $\mathsf{fdl}$$^{-}$ scores, hence improving explanation quality.